# Radial Profiles of Radio Halos in Massive Galaxy Clusters: Diffuse Giants Over 2 Mpc

### v1

@@ 5 Caveats on the profile extraction
The extraction of radial profiles is subject to systematic factors that can yield to misinterpretation of the results. In this section, we discuss key considerations that impact the interpretation of the halo radial profiles, including the subtraction/masking of discrete sources, the choice of sectors, and the selection of the sector center.
@@ 5.1 Source subtraction/masking
To assess the impact of removing unrelated radio sources from diffuse emission, we created profiles using two methods: (1) masking out all discrete sources (2) subtracting them from uv-data. The second method is commonly used approach in the literature.
@@
In the first method, discrete sources (compact) were masked based on the high and low frequency radio maps. We emphasize that to mask out extended unrelated sources, such as radio galaxies or relics embedded within the halo, we always used lowest frequency maps because these sources are typically more extended at low frequencies. We find that masking compact discrete sources using high and low frequency maps results in significantly different radial profiles. For example in Figures 11 and 8, we show the radial profiles of the Abell 2142 halo. Similar trends are observed in the radial profile of the other four halos analyzed (not shown). As evident when discrete sources are masked out based on the LOFAR HBA map (where the number density of compact sources is low), we see the presence of a second component at larger radius, see Figures 11 left. In contrast the profile obtained by masking out discrete sources based on the sensitive high frequency map, can be described by a single-component and shows no evidence of any second component (see Figure 8) .
@@
The secondary component is not real and results from the incomplete masking of discrete sources. In the inner region, the bright halo emission dominates. Therefore, incomplete source subtraction does not significantly impact the profiles. However, at larger radii, where the diffuse emission is weaker, the contribution from unmasked discrete sources becomes prominent relative to the halo emission, introducing an apparent shallower secondary component. This is shown in Figure 11 left panel where we present the residual profile (blue line) of the Abell 2142 halo, obtained by subtracting the radio profile with unrelated sources masked using the low-frequency map from the profile where masking was performed using the high-frequency map. As evident, in the outer regions ($>$600 kpc), where the halo surface brightness is low, the contribution from unmasked discrete sources becomes significant, resulting in an apparent second component at larger radial distances. Therefore, high frequency radio maps should be preferred for masking discrete sources as they typically show a higher density of point sources than low frequency maps due to better sensitivity.
@@
In the second method, we subtracted discrete (compact and/or extended) sources from the uv-data. In this method, we first created an image using an inner uv cut of $3.8\,k\lambda$ (with robust=$-0.5$), which was subsequently used as a model and subtracted from the uv data. The data were then re-imaged without any uv cut. For example in Figure 11 (right panel), we show the radial profile for the PLCK G287.0+32.9 halo at 815 MHz and uGMRT 350 MHz obtained by subtracting discrete from the uv-data. We note that regions covering the northern relic and the tailed radio galaxy to the northeast were manually masked. Unlike the profile obtained by directly masking out discrete sources from the image (Figure 7 top left), this profile shows a clear two-component structure at both 815 MHz and 350 MHz. The second component is is introduced due to incomplete subtraction of discrete sources.
@@
Figure 13: Top: Radial profiles extracted using circular, elliptical, and elliptical partial annuli. Profiles fitted with a single-component exponential model are shown with dashed lines. The left panel shows profiles extracted from an image where discrete unrelated sources within the halo were masked out, while the right panel shows profiles obtained from an image where discrete sources were subtracted from the $uv$-data, as commonly used in the literature.Bottom Left: Radial profiles in different sub-sectors; the dashed line shows the fit to the combined profile from all three sectors (Left). Bottom Right: Radial profiles extracted using different choices of centers. Bottom profiles were extracted from an image where discrete sources were subtracted from the $uv$-data. The plot demonstrates that the selection of sectors, sector centers, and discrete source subtraction significantly impacts the extracted radial profiles in radio halos with large extents and can introduce an artificial, shallower, second component at larger radii.
@@
Very recently, Salunkhe et al. (2025) reported uGMRT Band 3 and Band 4 observations of PLCK G287.0+32.9, claiming the presence of a mega-halo based on two components observed in the radial profile. However, our deeper observations reveal that the second component is not related to the halo. In Figure 12, we present 10″ resolution zoom-in images of the northern part of the halo using MeerKAT UHF and uGMRT Band 3 observations. We emphasize that our uGMRT Band 3 observation is a factor of two more sensitive than those reported by Salunkhe et al. (2025) at the same frequency range. Clearly, there are filamentary structures embedded within the halo region (marked with curved lines), as seen in our high-frequency, sensitive MeerKAT images. These features are not detected in the Band 3 images reported by Salunkhe et al. (2025) and are instead included as part of the halo emission. These structures appear more diffuse at lower frequencies. Moreover, the density of discrete sources is higher in our MeerKAT maps compared to the uGMRT Band3 map. As shown in the right panel of Figure 11, we also see a second, shallower component in both the 350 MHz and 815 MHz profiles when using an image where discrete sources are subtracted from the uv-data (based on the uGMRT Band3 image) as followed by Salunkhe et al. (2025). However, this second component is introduced by the incomplete subtraction of unrelated sources, in particular extended structures that are challenging to subtract from the uv-data. When discrete sources are properly masked out, based on sensitive high-frequency images, the radial profile of the PLCK G287.0+32.9 halo is well described by a single exponential fit (see Figure 7) at all the observed frequencies.
@@
Our analysis suggests that the subtraction/masking of discrete sources from the halo emission is a critical factor when characterizing radial profiles. We emphasize that we carried out this analysis using highly sensitive observations at high frequency and that the subtraction residuals would be exacerbated with images with poorer sensitivity to point sources and lower resolutions. We conclude that the masking process, which uses high-frequency data for compact sources and a combination of high and low frequency data for diffuse unrelated sources (e.g., filaments, relics, tailed radio galaxies) to exclude both thin structures recovered at high frequencies and emission with a steep spectrum detected at low frequencies, is the preferred approach for obtaining reliable radial profiles. If the discrete sources are primarily compact, subtracting them from the $uv$ data provides a reliable intrinsic halo radial profile. However, if extended sources are embedded within the halo emission, the masking approach should be preferred.
5.2 Sector choice
@@
Our analysis shows significant differences in the resulting profiles when using annuli with different geometries. In the top panels of Figure 13, we present the radial profile of the PLCK G287.0+32.9 halo, extracted from an image (with unrelated sources being masked out) using circular, elliptical, and elliptical partial annuli (see Figure 6 for the region descriptions). In the right panel, the same annuli are used, but with discrete sources were subtracted from the $uv$-data. When directly masking the discrete sources from the image, we observe only a marginal difference between the profiles obtained using different annuli (see Figure 13 left panel). However, a clear difference emerges when comparing these profiles to those extracted from a map where discrete sources were subtracted from the uv data, which shows the presence of a shallower second component (Figure 13 right panel). In the bottom left panel of Figure 13, we present additional profiles extracted from three sectors. These sectors are shown in Figure 6. Clearly, the choice of sector impacts the radial profiles. This is not surprising, as small sectors can capture substructures that may blend into smoothly varying emission when using circular or elliptical annuli. In conclusion, just focusing on a particular direction where the halo emission is enhanced may introduce a bias in the radial profiles.
5.3 Center of the sector
@@
To investigate how the selection of the center impacts the extracted profiles, we performed the analysis using multiple center positions and compared the resulting profiles, as shown in Figure 13 bottom right panel. As evident, the choice of the sector center introduces a noticeable difference in the profiles, particularly in the outer regions, where deviations become more pronounced.
@@
When the sector center is chosen at the peak of the radio emission, which is typically close to the X-ray peak, the resulting radial profile typically follows a single exponential fit. However, depending on the sector center, when multiple radio peaks are present in the core region (e.g., Abell 2142 and Bullet) or when discrete sources are embedded in the core region (e.g., PLCK,G287.0+32.9), the radial profile may show an apparent break at larger radii or deviate from a simple single-component structure. This effect becomes particularly significant when bright or extended discrete sources are located around the cluster core, as their subtraction can strongly affect the choice of the sector center, introducing biases into the radial profile analysis. For example, in PLCK G287.0+32.9, three bright compact sources are present in the core region, see Figure 2. Subtracting them from the $uv$-data could shifts the location of the radio peak, and thus the sector center, impacting the radial profiles (Figure 13, bottom right panel).
@@ 6 Conclusion
We analyzed the radial properties of five galaxy clusters, namely PLCK G287.0+32.9, Abell 2744, the Bullet Cluster, MACS J0717+3745, and Abell 2142 — all hosting radio halos. We find that despite their exceptionally large extents, the radio profiles of halos with LLS > 2 Mpc are well described by a single exponential fit, similar to those of classical radio halos. These five clusters are strong candidates for hosting mega-halos — of which four have been reported to date — based on their position in the cluster mass versus redshift plot and the total extent of their radio emission. We find no evidence of a second outer component in their radial profiles. While we were able to obtain profiles with a shallower second component in peripheral regions which is not real but rather a result of incomplete subtraction of discrete sources. We emphasize that we used high quality radio data to extract radial profiles. Our findings suggest that radio halos can extend to the cluster periphery, without the transition to an observationally distinguishable different halo component in the outermost regions.
@@
Despite their varying sizes, we find that the radio emissivity of halos with large extents is strikingly similar to that of classical halos, namely $10^{-42}\,\mathrm{erg\,s^{-1}\,cm^{-3}\,Hz^{-1}}$ (except Abell 2142). These halos with large extents lie above the known radio power at 1.4 GHz versus cluster mass scaling relations of halos (except Abell 2142). Moreover, their radial spectral index profiles reveal a clear spectral index gradient from the cluster core to the outer regions. Moreover, the powerful radio halos are found to be more extended, implying that such halos are expected to be detectable out to larger radii ($>$2 Mpc)
@@
Our findings demonstrate how the choice of sectors, sector centers, and discrete source subtraction considerably affect the radial profiles of radio halos. Careful subtraction of unrelated sources embedded in the halo is essential to ensure robust and reliable radial profiles. In particular, the identification of multiple components in radio halos with LLS > 2 Mpc based on radial profiles requires caution, as it can result in an artificial second, shallower component due to incomplete discrete source subtraction and the choice of sectors. Additional observational properties such as radial spectral index profiles and radio versus X-ray comparisons should be considered to assess the presence of a secondary component (e.g., Botteon et al., 2020; Rajpurohit et al., 2021a, b, 2022; Bruno et al., 2023; Biava et al., 2024).
@@
In conclusion, we find that halos with large extents (> 2 Mpc) share key characteristics with classical radio halos: 1) their averaged radial profiles are well described by a single-component exponential fit, 2) they show radial spectral index steepening, and 3) radio emissivity of $\sim 10^{-42}\,\mathrm{erg\,s^{-1}\,cm^{-3}\,Hz^{-1}}$. Our results highlight that the distinction between different categories of radio halos, such as hybrid halos, giant halos, and mega-halos, based on the size is becoming increasingly blurred. In contrast to the four mega halos with LLS> 2 Mpc, the five systems analyzed in this study do not exhibit different properties than those generally found in radio halos. Our results demonstrate that the observable size of a halo is primarily determined by the image depth and uv-coverage relative to its central brightness and e-folding radius. The $>2$ Mpc extents observed in some cases reflect the presence of bright, powerful radio halos, allowing their emission to be detected out to larger radii.
@@ Acknowledgments
KR acknowledges the Smithsonian Combined Support for Life on a Sustainable Planet, Science, and Research administered by the Office of the Under Secretary for Science and Research and NASA grant 619246. WF acknowledges support from the Smithsonian Institution, the Chandra High Resolution Camera Project through NASA contract NAS8-03060, and NASA Grants 80NSSC19K0116 and GO1-22132X. LB acknowledges support from the NextGenerationEU funds within the National Recovery and Resilience Plan (PNRR), Mission 4 - Education and Research, Component 2 -From Research to Business (M4C2), Investment Line 3.1- Strengthening and creation of Research Infrastructures, Project IR0000026- Next Generation Croce del Nord. The research leading to these results has received funding from the European Union?s Horizon 2020 research and innovation programme under grant agreement No 101004719 [ORP]. MeerKAT data reduction were conducted on the Smithsonian High Performance Cluster (SI/HPC), Smithsonian Institution https://doi.org/10.25572/SIHPC. The authors thank the staff of the MeerKAT observatory for their help with the observations presented in this work. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. We thank the staff of the GMRT that made these observations possible. GMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research. This research made use of the LOFAR-IT computing infrastructure supported and operated by INAF, including the resources within the PLEIADI special “LOFAR” project by USC-C of INAF, and by the Physics Dept. of Turin University (under the agreement with Consorzio Interuniversitario per la Fisica Spaziale) at the C3S Supercomputing Centre, Italy.

### v2

@@ 5 Caveats on the profile extraction
The extraction of radial profiles is subject to systematic factors that can lead to misinterpretation of the results. In this section, we discuss key considerations that impact the interpretation of the halo radial profiles, including the subtraction/masking of discrete sources, the choice of sectors, and the selection of the annulus/sector center.
@@ 5.1 Source subtraction/masking
To assess the impact of removing unrelated radio sources from halo emission, we created profiles using two methods: (1) masking out all discrete sources (2) subtracting them from uv-data using a uv-cut. The second method is a commonly used approach in the literature.
@@
In the first method, discrete sources (compact) were masked based on the high and low frequency radio maps. We emphasize that to mask out extended unrelated sources, such as radio galaxies or relics embedded within the halo, we always use the lowest frequency maps because these sources are typically more extended at low frequencies. We find that masking discrete sources using high and low frequency maps results in significantly different radial profiles. For example, in Figure 12 left panel, we show the radial profiles of the Abell 2142 halo obtained by masking sources based on the LOFAR 144 MHz and MeerKAT 1.3 GHz. Similar trends are observed in the radial profile of the other four halos analyzed. As evident when unrelated sources are masked out based on the LOFAR HBA map (where the number density of compact sources is low), we see the presence of a second component at larger radius, see Figures 12 left. In contrast, the profile obtained by masking out unrelated sources based on the sensitive high frequency map can be described by a single-component and shows no evidence of any second component.
@@
The secondary component is not real and results from the incomplete masking of unrelated sources. In the inner region, the bright halo emission dominates. Therefore, incomplete source subtraction does not significantly impact the profiles. However, at larger radii, where the diffuse emission is weaker, the contribution from unmasked discrete sources becomes prominent relative to the halo emission, introducing an apparent shallower secondary component. This is shown in Figure 12 left panel, where we present the residual profile (blue line) of the Abell 2142 halo, obtained by subtracting the radio profile with unrelated sources masked using the low-frequency map from the profile where masking was performed using the high-frequency map. As evident, in the outer regions (>>600 kpc), where the halo surface brightness is low, the contribution from unmasked discrete sources becomes significant, resulting in an apparent second component at larger radial distances. Therefore, high frequency radio maps should be preferred for masking discrete sources as they typically show a higher density of point sources than low frequency maps due to better sensitivity.
@@
In the second method, we subtracted discrete (compact and/or extended) sources from the uv-data. In this method, we first created an image using an inner uv-cut of $3.8\,k\lambda$ (with robust=0.5-0.5), which was subsequently used as a model and subtracted from the uv-data. We emphasize that the same uv-cut was applied to all sources, as this cut effectively recovered the flux from discrete sources. In none of the resulting maps was halo emission seen in the model image. The data were then re-imaged without any uv-cut. In Figure 7 last panel, we show the MeerKAT UHF (815 MHz) uv-subracted point source image. The resulting radial profiles for the PLCK G287.0+32.9 halo at 815 MHz and uGMRT 350 MHz is shown in Figure 12 right panel. The regions covering the northern relic and the tailed radio galaxy to the northeast were manually masked. We emphasize that in this case, the northernmost part of the halo is not masked and included in the radial profiles. Unlike the profile obtained by directly masking out discrete sources from the image (Figure 8 top left), this profile shows a two-component structure at both 815 MHz and 350 MHz. However, the second component is introduced due to the incomplete subtraction of discrete sources.
@@
Figure 15: Top: Radial profiles extracted using circular, elliptical, and elliptical partial annuli. Profiles fitted with a single-component exponential model are shown with dashed lines. For the used annuli see Figure 7. Middle: Radial profiles using different sectors; the dashed line shows the fit to the combined profile from all sectors. Sectors are shown in Figure 14 right panel. Bottom: Radial profiles extracted using different choices of centers. The left panels show profiles extracted from an image where unrelated sources within the halo were masked out, while the right panels show profiles obtained from an image where unrelated sources were subtracted from the uvuv-data, as commonly used in the literature. The plot demonstrates that the selection of sectors, sector centers, and discrete source subtraction significantly impacts the extracted radial profiles in radio halos with large extents, potentially introducing an artificial, shallower, second component at larger radii.
@@
Very recently, Salunkhe et al. (2025) reported uGMRT Band 3 and Band 4 observations of PLCK G287.0+32.9, claiming the presence of a mega-halo based on two components observed in the radial profile. However, our deeper observations reveal that the second component is not related to the halo. In Figure 13, we present 10″ resolution zoom-in images of the northern part of the halo using MeerKAT UHF and uGMRT Band 3 observations. We emphasize that our uGMRT Band 3 observation is a factor of two more sensitive than those reported by Salunkhe et al. (2025) at the same frequency range. Clearly, filamentary structures are embedded within the halo region (marked with curved lines), as seen in our high-frequency, sensitive MeerKAT images. These features are not detected in the Band 3 images reported by Salunkhe et al. (2025) and are instead included as part of the halo emission. These structures appear more diffuse at lower frequencies and are not associated with the halo emission but rather projected onto it (Rajpurohit et al. in prep). Moreover, the density of discrete sources is higher in our MeerKAT maps compared to the uGMRT Band3 map. As shown in the right panel of Figure 12, we also see a second, shallower component in both the 350 MHz and 815 MHz profiles when using an image where discrete sources are subtracted from the uv-data (based on the uGMRT Band3 image) as followed by Salunkhe et al. (2025). However, this second component is introduced by the incomplete subtraction of unrelated sources, in particular, extended structures that are challenging to subtract from the uv-data. When discrete sources are properly masked out, based on sensitive high-frequency images, the radial profile of the PLCK G287.0+32.9 halo is well described by a single exponential fit (see Figure 8) at all the observed frequencies.
@@
Our analysis suggests that the subtraction/masking of discrete sources from the halo emission is a critical factor when characterizing radial profiles. We emphasize that we carried out this analysis using highly sensitive observations at high frequency. We conclude that the masking process, which uses high-frequency data for compact sources and a combination of high and low frequency data for diffuse unrelated sources (e.g., filaments, relics, tailed radio galaxies) to exclude both thin structures recovered at high frequencies and emission with a steep spectrum detected at low frequencies, is the preferred approach for obtaining reliable radial profiles. If the discrete sources are primarily compact, subtracting them from the $uv$ data provides a reliable intrinsic halo radial profile. However, if extended sources are embedded within the halo emission, the masking approach should be preferred.
5.2 Annuli and Sector choice
@@
Our analysis shows significant differences in the resulting profiles when using annuli with different geometries. In the top-left panel of Figure 15, we present the radial profile of the PLCK G287.0+32.9 halo, extracted from an image (with unrelated sources being masked out) using circular, elliptical, and elliptical partial annuli (see Figure 7 and 14 left panel the annuli used) . In the top right panel, the same annuli were used to extract the surface brightness but from an image where unrelated sources were subtracted from the $uv$-data. When directly masking the unrelated sources from the image, we observe only a marginal difference between the shape of the profiles obtained using different annuli and there is no evidence of the second shallower component at larger radii (see Figure 15 top-left panel). However, a clear difference emerges when comparing these profiles to those extracted from a map where unrelated sources were subtracted from the uv data, which shows the presence of a shallower second component (Figure 15 top-right panel). It is worth noting that in Figure 15 top right panel profile, the northernmost part of the halo is also included. In the middle panels of Figure 15, we present additional profiles extracted from different sectors. These sectors are shown in Figure 14 left panel. Clearly, the choice of sector impacts the radial profiles regardless of whether a masked or uv-subtracted image is used. This is not surprising, as small sectors can capture substructures that may blend into smoothly varying emission or regions with different properties when using different sectors. In conclusion, just focusing on a particular direction where the halo emission is enhanced may introduce a bias in the radial profiles.
5.3 Center of the annuli
@@
To investigate how the selection of the center impacts the extracted profiles, we performed the analysis using multiple center positions and compared the resulting profiles, as shown in Figure 15 bottom panels. The used center positions are shown in Figure 14. right panel. As evident, the choice of the annuli center introduces a noticeable difference in the profiles, particularly in the outer regions, where deviations become more pronounced.
@@
When the sector center is chosen at the peak of the radio emission, which is typically close to the X-ray peak, the resulting radial profile typically follows a single exponential fit. However, depending on the sector center, when multiple radio peaks are present in the core region (e.g., Abell 2142 and Bullet) or when discrete sources are embedded in the core region (e.g., PLCK G287.0+32.9), the radial profile may show an apparent break at larger radii or deviate from a simple single-component structure. This effect becomes particularly significant when bright or extended discrete sources are located around the cluster core, as their subtraction can strongly affect the choice of the sector center, introducing biases into the radial profile analysis. For example, in PLCK G287.0+32.9, three bright compact sources are present in the core region, see Figure 2. Subtracting them from the $uv$-data could shift the location of the radio peak, and thus the sector center, impacting the radial profiles (Figure 15, bottom panels).
@@ 6 Conclusion
We analyzed the radial properties of five galaxy clusters, namely PLCK G287.0+32.9, Abell 2744, the Bullet Cluster, MACS J0717+3745, and Abell 2142 — all hosting radio halos. We find that despite their exceptionally large extents, the radio profiles of halos with LLS > 2 Mpc can be described by a single exponential fit, similar to those of classical radio halos. These five clusters are strong candidates for hosting mega-halos — of which four have been reported to date — based on their position in the cluster mass versus redshift plot and the total extent of the halo emission. We find no evidence of a second outer component in their radial profiles. While we were able to obtain profiles with a shallower second component in peripheral regions, it is not real but rather a result of incomplete subtraction of unrelated sources. We emphasize that we used high quality radio data to extract radial profiles. Our findings suggest that radio halos can extend to the cluster periphery, without the transition to an observationally distinguishable different halo component in the outermost regions.
@@
Despite their varying sizes, we find that the radio emissivity of halos with large extents is strikingly similar to that of classical halos, namely $10^{-42}\,\mathrm{erg\,s^{-1}\,cm^{-3}\,Hz^{-1}}$ (except Abell 2142). These halos with large extents lie above the known radio power at 1.4 GHz versus cluster mass scaling relations of halos (except Abell 2142). Moreover, their radial spectral index profiles reveal a clear spectral index gradient from the cluster core to the outer regions. Moreover, the powerful radio halos are found to be more extended, implying that such halos are expected to be detectable out to larger radii (>>2 Mpc)
@@
Our findings demonstrate that the choice of sectors, annulus centers, and discrete source subtraction significantly affects the radial profiles of halos. Careful subtraction of unrelated sources embedded in the halo is essential to ensure robust and reliable radial profiles. In particular, the identification of multiple components in radio halos based on radial profiles requires caution, as it can result in an artificial second, shallower component due to incomplete subtraction of unrelated sources and the choice of annuli/sectors. Additional observational properties such as radial spectral index profiles and radio versus X-ray comparisons should be considered to assess the presence of a secondary component (e.g., Botteon et al., 2020; Rajpurohit et al., 2021a, b, 2022b; Bruno et al., 2023; Biava et al., 2024).
@@
In conclusion, we find that halos with large extents (>2>2 Mpc) share key characteristics with classical radio halos: 1) their averaged radial profiles are well described by a single-component exponential fit, 2) they show radial spectral index steepening, and 3) radio emissivity of $\sim 10^{-42}\,\mathrm{erg\,s^{-1}\,cm^{-3}\,Hz^{-1}}$. Our results highlight that the distinction between different categories of radio halos, such as hybrid halos, giant halos, and mega-halos, based on size is becoming increasingly blurred. In contrast to the four mega halos with $\rm LLS>2\,Mpc$, the five systems analyzed in this study do not exhibit different properties than those generally found in radio halos. Our results demonstrate that the observable size of a halo is primarily determined by the image depth and uv-coverage relative to its central brightness and e-folding radius. The >2>2 Mpc extents observed in some cases reflect the presence of bright, powerful radio halos, allowing their emission to be detected out to larger radii.
@@ Acknowledgments
KR acknowledges funding support from the Smithsonian Combined Support for Life on a Sustainable Planet, Science, and Research administered by the Office of the Under Secretary for Science and Research and NASA grant 619246. WF acknowledges support from the Smithsonian Institution, the Chandra High Resolution Camera Project through NASA contract NAS8-03060, and NASA Grants 80NSSC19K0116 and GO1-22132X. CJ acknowledges support from the Smithsonian Institution. LB acknowledges support from the NextGenerationEU funds within the National Recovery and Resilience Plan (PNRR), Mission 4 - Education and Research, Component 2 -From Research to Business (M4C2), Investment Line 3.1- Strengthening and creation of Research Infrastructures, Project IR0000026- Next Generation Croce del Nord. The research leading to these results has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101004719 [ORP]. MeerKAT data reduction were conducted on the Smithsonian High Performance Cluster (SI/HPC), Smithsonian Institution https://doi.org/10.25572/SIHPC. The authors thank the staff of the MeerKAT observatory for their help with the observations presented in this work. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. We thank the staff of the GMRT that made these observations possible. GMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research. This research made use of the LOFAR-IT computing infrastructure supported and operated by INAF, including the resources within the PLEIADI special “LOFAR” project by USC-C of INAF, and by the Physics Dept. of Turin University (under the agreement with Consorzio Interuniversitario per la Fisica Spaziale) at the C3S Supercomputing Centre, Italy.